# Large Language Models are In-Context Molecule Learners

### v2

@@ A.3.1 Results with More Backbone LLMs
To highlight the model agnosticism of ICMA, we have expanded our experiments on two more backbone LLMs, Galactica-1.3B and Llama-2-7b-chat-hf. The results are shown in Tables 11 and 12. It is evident that ICMA enhances the performance of both models, which further proves the versatility of ICMA.
@@
Llama2-7B 0.515 0.423 0.583 0.412 0.518 0.528
ICMA(Llama2-7B) 0.629 0.553 0.663 0.518 0.601 0.635
Table 11: Performance comparison of more backbone LLMs for Mol2Cap task on ChEBI-20 dataset. Here, we select Galactica-1.3B and Llama-2-7b-chat-hf for demonstration.
@@
Llama2-7B 0.663 0.130 37.24 0.801 0.649 0.572 0.861
ICMA(Llama2-7B) 0.783 0.319 27.74 0.888 0.792 0.727 0.923
Table 12: Performance comparison of more backbone LLMs for Cap2Mol task on ChEBI-20 dataset. Here, we select Galactica-1.3B and Llama-2-7b-chat-hf for demonstration.
A.3.2 Comparison with Other Baseline Models
In our pursuit to comprehensively evaluate the efficacy of ICMA, we have extended our comparative analysis to include additional baseline models. Specifically, we first compare ICMA(Mistral-7B) with BioT5  Pei et al. (2023),  Liu et al. (2023a), and MolCA  Liu et al. (2023b) in the Mol2Cap task. The results, as detailed in Table 13, indicate that ICMA outperforms all the baseline models in terms of BLEU-2, BLEU-4, and METEOR scores, while also achieving the second-highest ROUGE scores, which further prove that ICMA could empower Mistral-7B to achieve SOTA or comparable performance.
Notably, compared to MolCA, a multi-modal method that leverages 2D molecular graphs to augment LLMs for the Mol2Cap task, our ICMA is capable of refining the alignment between molecular and textual representations not only for Mol2Cap task but also for Cap2Mol task, which is more important to the field of molecule discovery.
Furthermore, for the Cap2Mol task, we have included BioT5  Pei et al. (2023) and MolXPT  Liu et al. (2023a) in our comparative analysis. As presented in Table 14, ICMA achieves the highest scores in exact match and molecule fingerprints metrics, indicative of superior quality in molecule generation. However, it is also important to note that since the Mistral-7B model was not pre-trained on a large-scale chemical corpus, the validity of the generation is slightly below the two baseline models due to a deficiency in molecular knowledge.
@@ A.3.2 Comparison with Other Baseline Models
BioT5  Pei et al. (2023) 0.635 0.556 0.692 0.559 0.633 0.656
MolXPT  Liu et al. (2023a) 0.594 0.505 0.660 0.511 0.597 0.626
MolCA  Liu et al. (2023b) 0.620 0.531 0.681 0.537 0.618 0.651
@@
Table 13: Performance comparison with more baseline models for Mol2Cap task on ChEBI-20 dataset. Here, we select BioT5  Pei et al. (2023), MolXPT  Liu et al. (2023a), and MolCA  Liu et al. (2023b) for comparison.
@@
Table 14: Performance comparison of more baseline models for Cap2Mol task on ChEBI-20 dataset. Here, we include the results from BioT5  Pei et al. (2023) and MolXPT  Liu et al. (2023a) for comparison.
A.4 Case Study
@@
Mol2Cap Figure 6 presents a set of caption examples produced by two baseline methods and our ICMA. In the first example, MolT5 incorrectly equated D-tartrate(2-) with L-tartrate(2-) and L-tartrate(1-) with D-tartrate(1-), while MolReGPT introduced an excessive amount of irrelevant information. In contrast, ICMA could accurately empower backbone models to generate precise answers. Meanwhile, in the second example, ICMA accurately recognized that the molecule in question "has a role as an antibacterial drug", whereas MolReGPT failed. Moreover, in the fourth example, ICMA meticulously captured all the essential details in describing the molecule structures without a single mistake, demonstrating that contextual examples can significantly enhance the understanding of molecule structures, especially the types and positions of functional groups.
To summarize, ICMA could fine-grain the detailed alignment between molecule captions and molecule SMILES representations via learning from context examples. ICMA could match the types and positions of functional groups, the number of carbon atoms, as well as the overall structures more accurately, showing that LLMs are inherently in-context molecule learners.

### v3

@@ IV-F Study of Model Agnosticism
To highlight the model agnosticism of ICMA, we also expand our experiments on two more backbone LLMs, Galactica-1.3B and Meta-Llama-3-8B-Instruct. The results are shown in Tables IX and X. It is evident that ICMA enhances the performance of both backbone models, which further proves the versatility and model agnosticism of ICMA.
TABLE IX: Performance comparison of two more backbone LLMs (Galactica-1.3B and Meta-Llama-3-8B-Instruct) for Mol2Cap task on ChEBI-20 dataset (Best, Second Best).
@@
Llama3-8B 0.618 0.542 0.660 0.515 0.599 0.623
ICMA(Llama3-8B) 0.665 0.595 0.693 0.559 0.633 0.669
TABLE X: Performance comparison of two more backbone LLMs (Galactica-1.3B and Meta-Llama-3-8B-Instruct) for Cap2Mol task on ChEBI-20 dataset (Best, Second Best).
@@
Llama3-8B 0.826 0.356 20.70 0.894 0.793 0.733 0.939
ICMA(Llama3-8B) 0.851 0.445 19.27 0.915 0.836 0.785 0.958
IV-G Performance for Complex Molecules and Captions
ICMA also works better for complex molecular structures and molecule captions. Normally, the lengthy molecules can be more complex for LLMs. In this case, we extracted a subset of 747 complex molecules, whose SMILES strings have more than 100 characters and a subset of 221 complex captions with more than 500 characters from the ChEBI-20 test set. We compared the performance of ICMA with previous baselines on this subset. Our experimental results shown in Table XI and XII indicate that ICMA outperforms the previous baselines, especially in the Cap2Mol task, as similar examples learned from ICMT typically share comparable levels of complexity, which helps better aligns molecules with texts.
TABLE XI: The Mol2Cap performance comparison on the subset of ChEBI-20 with complex molecules whose SMILES strings have more than 100 characters (Best, Second Best).
@@ IV-G Performance for Complex Molecules and Captions
MolT5-large 0.684 0.624 0.715 0.590 0.661 0.685
MolReGPT(GPT-4-0314) 0.670 0.603 0.692 0.555 0.630 0.673
ICMA(Mistral-7B) 0.687 0.632 0.727 0.611 0.673 0.713
TABLE XII: The Cap2Mol performance comparison on the subset of ChEBI-20 with complex captions that have more than 500 characters (Best, Second Best).
[…]
MolT5-large 0.624 0.113 50.71 0.801 0.669 0.584 0.792
MolReGPT(GPT-4-0314) 0.794 0.122 40.00 0.858 0.704 0.630 0.769
ICMA(Mistral-7B) 0.684 0.235 70.79 0.883 0.736 0.670 0.873
IV-H Random Walk Stability & Maximum Skip Probability
Due to the randomness we introduced during the Random Walk selection, there might be concerns about the stability of the strategy. However, considering that the design of the skip probability is to limit it within the maximum skip probability, which is a rather low value, and gradually decays the likelihood to ensure that at least one example is selected, we manage the randomness brought by Random Walk selection within an acceptable range. To verify this, we implemented a bucket sampling strategy for comparison. We randomly chose three different random seeds and calculated the mean and variance of the performance results for three runs. The results are demonstrated in Table XIII and XIV.
TABLE XIII: Stability Comparison on the Mol2Cap task between two re-ranking strategy, Random Walk and Bucket Sampling. The results are with significant figures for precision (Best).
[…]
Random Walk 0.6392±0.0032 0.5658±0.0031 0.6703±0.0028 0.5306±0.0026 0.6107±0.0030 0.6437±0.0031
Bucket Sampling 0.6364±0.0035 0.5636±0.0040 0.6687±0.0032 0.5293±0.0033 0.6087±0.0026 0.6424±0.0043
TABLE XIV: Stability Comparison on the Cap2Mol task between two re-ranking strategy, Random Walk and Bucket Sampling. The results are with four significant figures for precision (Best).
[…]
Random Walk 0.8403±0.0024 0.3533±0.0026 18.45±0.56 0.8907±0.0020 0.8042±0.0033 0.7412±0.0031 0.9336±0.0027
Bucket Sampling 0.8371±0.0014 0.3464±0.0063 18.58±0.23 0.8885±0.0020 0.8000±0.0021 0.7367±0.0021 0.9347±0.0056
From the results, we can observe that although the variance of the Random Walk is larger than that of bucket sampling on metrics like BLEU, Levenshtein, RDK FTS, and Morgan FTS in the Cap2Mol task and ROUGE-L in the Mol2Cap task, the Random Walk strategy still achieves higher means, and the variances are still manageable. On some metrics, the variances of the Random Walk are even lower than that of bucket sampling, indicating that despite the inherent randomness, our method overall remains stable within normal fluctuation ranges.
Figure 5: The performance of Galatcica-125M with the maximum skip probability $p_{max}$. Mol2Cap Results (Left) and Cap2Mol Results (Right).
What’s more, we further conduct a hyperparamter search to help decide the best maximum skip probability $p_{max}$ under the setting of $n=2$, $N=10$, and $cuto\!f\!f\_len=1024$. As shown in Figure 5, we could obviously see that if the maximum skip probability is too large, the performance naturally drops because it is more likely to select less similar examples. However, with the assistance of skip probability decay as well as the early-stop, the performance gap is still marginal. Among all the selected maximum probabilities, $p_{max}=9\%$ achieves the highest exact match score in the Cap2Mol task and the best METEOR score in the Mol2Cap task. However, we must acknowledge that the Random Walk strategy inherently introduces a degree of randomness, and for different numbers of rough examples $N$ and refined examples $n$, the best maximum probability might vary, which means the observed results may not fully reflect the true potential of the Random Walk sampling.
IV-I Ablation Study
TABLE XV: Ablating components of Post-retrieval Re-ranking for Mol2Cap task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
[…]
ICMA 0.641 0.568 0.671 0.531 0.611 0.645
w/o Random Walk 0.638(3) 0.565(5) 0.670 0.530(3) 0.610(3) 0.643(7)
w/o Sequence Reverse 0.638(4) 0.566(3) 0.669 0.530(5) 0.610(1) 0.644(1)
TABLE XVI: Ablating components of Post-retrieval Re-ranking for Cap2Mol task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
[…]
ICMA 0.842 0.355 17.99 0.890 0.805 0.741 0.935
w/o Random Walk 0.840 0.348 18.01 0.888 0.804 0.738 0.939
w/o Sequence Reverse 0.835 0.354 18.49 0.889 0.800 0.740 0.933
Ablation study is also conducted to illustrate how the Post-retrieval Re-ranking components affect the predictions. We conduct ablation experiments for the Mol2Cap and Cap2Mol sub-tasks by deactivating the Random Walk and Sequence Reversal. As shown in Table XV and XVI, without Random Walk or Sequence Reversal, the performance of ICMA drops, though not significantly but consistently, in both tasks, proving the effectiveness of Post-retrieval Re-ranking.
IV-J Comparative Analysis of Models with Extra Domain Alignment Stages
In our pursuit to comprehensively evaluate the efficacy of ICMA, we have extended our comparative analysis to include models that adopt extra domain alignment stages, such as continual pre-training on chemical corpora (BioT5 [12] & MolXPT [11]) and graph modality alignment (MolCA  [9]. The results, as detailed in Table XVII, indicate that ICMA outperforms all the models trained with extra domain alignment in terms of BLEU-2, BLEU-4, and METEOR scores, while also achieving the second-highest ROUGE scores. Notably, the performance is achieved without introducing extra domain alignment stages, which further proves the superiority of ICMA.
Notably, compared to MolCA, a multi-modal method that leverages 2D molecular graphs to augment LLMs for the Mol2Cap task, our ICMA does not require extra graph information and modality alignment training. Meanwhile, ICMA is capable of refining the alignment between molecular and textual representations not only for Mol2Cap task but also for Cap2Mol task, showing better versatility in aligning molecular space with textual space.
On the other hand, for the Cap2Mol task, we have included BioT5 and MolXPT in our comparative analysis. As presented in Table XVIII, ICMA achieves the highest scores in exact match and molecule fingerprints metrics, while the validity of molecule generation is just slightly below the BioT5 and MolXPT due to the deficiency in molecular knowledge, indicating competitive performance in molecule generation. Furthermore, it is also important to note that ICMA does not require pre-training LLMs on a large-scale chemical corpus, which makes ICMA a better choice in low-data scenarios.
TABLE XVII: Performance comparison with models trained with extra domain alignment stages for Mol2Cap task on ChEBI-20 dataset. Here, we select BioT5 [12], MolXPT [11], and MolCA [9] for comparison (Best, Second Best).
[…]
BioT5 [12] 0.635 0.556 0.692 0.559 0.633 0.656
MolXPT [11] 0.594 0.505 0.660 0.511 0.597 0.626
MolCA [9] 0.620 0.531 0.681 0.537 0.618 0.651
@@ IV-J Comparative Analysis of Models with Extra Domain Alignment Stages
TABLE XVIII: Performance comparison with models trained with extra domain alignment stages for Cap2Mol task on ChEBI-20 dataset. Here, we include the results from BioT5 [12] and MolXPT [11] for comparison (Best, Second Best).
@@
IV-K ICMA for Molecule Property Prediction
Although ICMA is targeted at the task of molecule-caption translation, we are curious about whether ICMA could also benefit a wide range of molecule-related tasks, such as molecule property prediction. Therefore, we conduct experiments on the MoleculeNet dataset [31] and mainly focus on classification subtasks like BACE, BBBP, HIV, TOX21, and SIDER. The results are shown in Table XIX. To our surprise, ICMA shows great generalization capability to these molecule property prediction tasks, increasing the performance of both Galactica-125M and Mistral-7B. More significantly, ICMA enables the general LLM, Mistral-7B, to achieve an average improvement of 33.26% on the five subtasks, which further proves that LLMs are inherently in-context molecule learners.
TABLE XIX: Molecule Property Prediction Performance of ICMA compared with directly fine-tuning the backbone models on the MoleculeNet Dataset (Best, Second Best).
[…]
Galactica-125M 0.8216 0.7805 0.7314 0.7453 0.7642
ICMA (Galactica-125M) 0.8941 0.8680 0.7412 0.7476 0.7673
Mistral-7B 0.4926 0.4829 0.5866 0.4386 0.3771
ICMA (Mistral-7B) 0.7995 0.6775 0.7441 0.4761 0.4838
IV-L Case Study
@@
Mol2Cap Figure 7 presents a set of caption examples produced by two baseline methods and our ICMA. In the first example, MolT5 incorrectly equated D-tartrate(2-) with L-tartrate(2-) and L-tartrate(1-) with D-tartrate(1-), while MolReGPT introduced an excessive amount of irrelevant information. In contrast, ICMA could accurately empower backbone models to generate precise answers. Meanwhile, in the second example, ICMA accurately recognized that the molecule in question ”has a role as an antibacterial drug”, whereas MolReGPT failed. Moreover, in the fourth example, ICMA meticulously captured all the essential details in describing the molecule structures without a single mistake, demonstrating that contextual examples can significantly enhance the understanding of molecule structures, especially the types and positions of functional groups.
Overall, ICMA could fine-grain the detailed alignment between molecule captions and molecule SMILES representations via learning from context examples, showing better capabilities in matching the types and positions of functional groups, the number of carbon atoms, as well as the overall structures.
@@
V Conclusion
In this work, we propose In-Context Molecule Adaptation (ICMA), as a new paradigm for adapting LLMs to the molecule-caption translation task. Instead of domain-specific pre-training and fine-tuning, ICMA enables LLMs to utilize their in-context learning capability to learn molecule-text alignment via In-Context Molecule Tuning, which significantly improves the performance of LLMs in the molecule-caption translation task, demonstrating that LLMs are inherently in-context molecule learners. More importantly, our study provides a viable framework for deploying advanced LLMs with billion-level parameters without extra domain alignment stages in the scientific field, making it suitable to be applied in low-data scenarios.
VI Acknowledgement
The research described in this paper has been partly supported by the General Research Funds from the Hong Kong Research Grants Council (project no. PolyU 15200023, and 15224524). This work is also supported by Shanghai Artificial Intelligence Laboratory (Shanghai AI Lab) and was done during Jiatong Li and Wei Liu’s internship at Shanghai AI Lab.